OpenStudy (anonymous):

Solve for x: sin^2x = 4-2cos^2x

OpenStudy (solomonzelman):

\(\large\color{black}{ \sin^2x = 4-2\cos^2x }\) there is a rule, \(\large\color{red}{ \sin^2x+\cos^2x=1 }\) and when we subtract \(\large\color{red}{ \cos^2x }\) from both sides, we get \(\large\color{red}{ \sin^2x=1 -\cos^2x }\) So, \(\large\color{black}{ (1-\cos^2x )= 4-2\cos^2x }\)

OpenStudy (solomonzelman):

(have re-written sin^2x in terms of cos^2x. )

OpenStudy (solomonzelman):

Tell me if you need more help from this point and on.

OpenStudy (anonymous):

so the cos should be sqrt(3)?

OpenStudy (solomonzelman):

Let's see.... \(\large\color{black}{ 1-\cos^2x =4-2\cos^2x }\) adding cos^2x to both sides, \(\large\color{black}{ 1 =4-3\cos^2x }\) subtract 4 from both sides \(\large\color{black}{ -3 =-3\cos^2x }\)

OpenStudy (solomonzelman):

can you take it from here?

OpenStudy (anonymous):

i got \[\cos ^{2} x = 3\] therefore \[\cos x = \sqrt{3}\]

OpenStudy (solomonzelman):

after the last thing I wrote, I proceed by dividing by -3 on both sides, I get, \(\large\color{black}{ \cos^2x=1 }\) then I take a square root of both sides, \(\large\color{black}{ \cos x=\pm1 }\) \(\large\color{blue }{ x=1,~~\pi }\)

OpenStudy (anonymous):

but -2cos^2x + cos^2 x = -cos^2 x

OpenStudy (solomonzelman):

yes....

OpenStudy (solomonzelman):

and so?

OpenStudy (dumbcow):

lets see i think you are correct @KRJones23

OpenStudy (dumbcow):

@SolomonZelman , when you added cos^2 .... it should have been -2+1 = -1

OpenStudy (solomonzelman):

ohh, true lol.

OpenStudy (solomonzelman):

so then 3=cos^2x.

OpenStudy (dumbcow):

cos x must be within -1 and 1 so this equation has no solution

OpenStudy (solomonzelman):

yuup. (shouldn't be saying that, because sounds like I am the one who "knows", but still yes)